OpenStudy (anonymous):

What kind of equation is 5x-4y=8?

OpenStudy (shubhamsrg):

0 degree linear differential eqn..

OpenStudy (anonymous):

Thank you!

OpenStudy (shubhamsrg):

:D

OpenStudy (shubhamsrg):

what's the topic of your homework ? am assuming 2-d,,or more specifically, straight lines ?

OpenStudy (shubhamsrg):

bah alright, i was bluffing , though thats correct,,still,that not the ans you were looking for..

OpenStudy (aravindg):

its a linear equation

OpenStudy (anonymous):

my homework problem is to find the slope-intercept forms of the equations of the line s through the given point a. parallel to the given line b. perpendicular to the given line. line is 5x-4y=8 and point is (3,2)

OpenStudy (shubhamsrg):

read your last line again and you shall know what kind of eqn is 5x-4y=8

OpenStudy (anonymous):

a perpendicular one?

OpenStudy (anonymous):

im beyond help man.. They shouldn't have let me into Pre-Calculus. I'm stumped on the freakin review!

OpenStudy (tyteen4a03):

Hint: The slope-intercept form of a line is \(y = mx + c\), where m is the slop and c the y intercept. Parallel lines have the same slope. Perpendicular lines have negative reciprocals slopes. (i.e 2 VS -1/2) For this question, find the y-intercept by plugging in that x = 0 (y intercept lies on the origin, hence x will be zero), then find the slope of (0, y-intercept) and (3,2). The slope formula is \(\frac{y2 - y1} {x2 - x1}\). Then, for a, just add some random numbers to c to get a parallel line. For b, simply replace m with the negative reciprocals of m.